# 1.3in1.3in1in1in

### v1

@@ 5. Appendix
Under a suitable change of coordinate every cubic surface with 4 singular points of type $A_{1}$ can be expressed as zero locus of the following polynomial:
$F=t_{0}t_{1}t_{2}+t_{0}t_{1}t_{3}+t_{0}t_{2}t_{3}+t_{1}t_{2}t_{3}$
@@
Such a surface can be obtained as the anticanonical model of the blow-up of $\mathbb{P}^{2}$ in special configuration of points. Indeed, let $\widetilde{S}$ be the blow-up of $P^{2}$ in the six points given by the intersection of 4 general lines $L_{0},L_{1},...,L_{3}$. Then the anticanonical linear system of $\widetilde{S}$ maps to a cubic surfaces $S$ and contracts the strict tranforms of $L_{i}$ to 4 singular $A_{1}$ points. The exceptional divisors map to lines in a tetrahedron-like disposition, we refer to them as edges, and denote $E_{ij}$ the line connecting the $i$-th and the $j$-th coordinate point. In particular, the edge $E_{ij}$ is cut by the equations $(t_{k}=t_{l}=0)$ where $\{i,j,k,l\}=\{1,2,3,4\}$. Besides the edges there are 3 more lines on $\bar{S}$ Joining opposite edges; we denote by $J_{ij,kl}$ the line intersecting the edges $E_{ij}$ and $E_{kl}$, where $\{i,j,k,l\}=\{1,2,3,4\}$. These are images of the strict transforms of lines through the blown-up points corresponding to the edges it intersects. For example, the line $J_{01,23}$ is the zero locus $(t_{0}+t_{1}=t_{2}+t_{3}=0)$.
@@
Lemma 5.1.
Opposite edges map to $K$ under $\varphi$, that is:
@@
The proof is elementary and can be checked directly by the definition of the Voisin map $\varphi$. For example to compute $\varphi(E_{02},E_{13})$, consider the plane $\pi=(t_{1}=0)$ containing the line $E_{02}$ and the point $p_{\in}E_{13}$. Its intersection with $S$ is given by $\pi\cap S=E_{02}\cup E_{03}\cup E_{23}$, hence $\varphi(E_{02},E_{13})=\left[E_{03}\cup E_{13}\cup E_{23}\right]$ and the claim is shown.
@@
Let $S$ be a cubic surface with a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{01}$ through $P_{0},P_{1}$; the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{03}$ through $P_{03}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
@@
The lines are then described by the equations $R_{23}=\{t_{0}=t_{1}=0\},R_{02}=\{t_{1}=t_{3}=0\}$ and $R_{03}=\{t_{1}=t_{2}=0\}$, $L_{1}=(t_{0}=t_{3}=0)$, and $L_{2}=(t_{2}=t_{0}+t_{3}=0)$.
Proposition 5.2.
Let $S$ be a cubic surface of type $2A_{1}A_{3}$. Then:
$L_{1}\mbox{ is residual to }R_{23},\qquad L_{1}\mbox{ is residual to }R_{02},$
$L_{2}\mbox{ is residual to }R_{03},\qquad L_{2}\mbox{ is residual to }L_{1}.$
Moreover, there is no plane in $\langle S\rangle\simeq\mathbb{P}^{3}$ tangent to $S$ at every point of $L_{2}$. In particular, there are no triple lines on a surface with $2A_{1}A_{3}$ singularities.
@@
The proof can be checked through straightforward computations. For example, intersecting the surface $S$ with the plane $\pi=(t_{0}=0)$, one gets:
@@
whose singular locus consists of the 2 points $P_{1}:=(0:0:0:1)$ of type $A_{5}$ and $P_{2}:=(0:0:1:0)$ of type $A_{1}$. There are exactly 2 lines on $S$ and they have equations $L_{1}=\{t_{0}=t_{1}=0\}$ and $L_{2}=\{t_{0}=t_{3}=0\}$.
Proposition 5.3.
Let $S$ be a cubic surface of type $A_{1}A_{5}$. Then
$L_{2}\text{ is triple, and }L_{2}\text{ is residual to }L_{1}.$
@@
Straightforward computation: The intersection
$\{t_{3}=0\}\cap S=3L_{2}.$
shows the first statement, while
$\{t_{0}=0\}\cap S=L_{1}+2L_{2}$
shows the second one. ∎
@@
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.4.
@@
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{0}+bt_{1}=0)$ with $(a:b)\in\mathbb{P}^{1}$. By considering the intersection with such a plane $\pi$ one has:
$X_{6}\cap(-t_{0}+bt_{1}=0)=(-t_{0}+bt_{1}=t_{1}^{2}(b^{2}t_{2}+t_{3})=0)=2L+R$
with $R=b^{2}t_{2}+t_{3}=bt_{1}+t_{0}=0$ different from $L$, and
$X_{6}\cap(t_{1}=0)=(t_{1}=t_{0}^{2}t_{2}=0)=2L+R$
@@
$t_{1}^{3}+t_{2}^{3}+t_{1}t_{2}t_{3}.$
It is a non-normal surface isomorphic to the cone over a nodal planar cubic whose singular locus consists of the line $L=(t_{1}=t_{2}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.5.
@@
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{1}+bt_{2}=0)$ with $(a:b)\in\mathbb{P}^{1}$.
By considering the intersection with such a plane $\pi$ one has:
$X_{7}\cap(t_{0}-bt_{1}=0)=(t_{0}-bt_{1}=t_{1}^{2}(bt_{2}+b^{2}t_{3}+t_{1})=0)=2L+R.$
with $R=(t_{0}-bt_{1}=t_{1}+bt_{2}+b^{2}t_{3}=0)$. Also, one has:
$X_{7}\cap(t_{0}=0)=(t_{0}=t_{1}^{3}=0)=3L.$
This shows the claim. ∎
@@
$t_{0}t_{1}t_{2}+t_{0}^{2}t_{3}+t_{1}^{3}=0.$
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$. We record only the following which is necessary for the proof of Corollary 3.20
Proposition 5.6.
@@
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi=(at_{1}+bt_{2}=0)$ with $(a:b)\in\mathbb{P}^{1}$.
By considering the intersection with such a plane $\pi$ one has:
$X_{8}\cap(t_{1}-bt_{2}=0)=(t_{1}-bt_{2}=t_{2}^{2}(b^{3}t_{2}+t_{2}+bt_{3})=0)=2L+R.$
with $R=(t_{0}-bt_{1}=t_{1}+bt_{2}+b^{2}t_{3}=0)$. Also, one has:
$X_{8}\cap(t_{2}=0)=(t_{1}^{3}=t_{2}=0)=3L.$

### v2

@@ Appendix A Singular cubic surfaces
Under a suitable change of coordinate every cubic surface with 4 singular points of type $A_{1}$ can be expressed as zero locus of the polynomial
$t_{0}t_{1}t_{2}+t_{0}t_{1}t_{3}+t_{0}t_{2}t_{3}+t_{1}t_{2}t_{3}.$
@@
Such a surface can be obtained as the anticanonical model of the blow-up of $\mathbb{P}^{2}$ in special configuration of points. Indeed, let $\widetilde{S}$ be the blow-up of $\mathbb{P}^{2}$ in the six points given by the intersection of 4 general lines $L_{0},L_{1},...,L_{3}$. Then the anticanonical linear system of $\widetilde{S}$ maps to a cubic surfaces $S$ and contracts the strict transforms of $L_{i}$ to 4 singular $A_{1}$ points. The exceptional divisors map to lines in a tetrahedron-like arrangement, we refer to them as edges, and denote $E_{ij}$ the line connecting the $i$-th and the $j$-th coordinate point. In particular, the edge $E_{ij}$ is cut by the equations $(t_{k}=t_{l}=0)$ where $\{i,j,k,l\}=\{0,1,2,3\}$. Besides the edges there are 3 more lines on $S$ Joining opposite edges; we denote by $J_{ij,kl}$ the line intersecting the edges $E_{ij}$ and $E_{kl}$, where $\{i,j,k,l\}=\{0,1,2,3\}$. These are images of the strict transforms of lines through the blown-up points corresponding to the edges it intersects. For example, the line $J_{01,23}$ is the zero locus $(t_{0}+t_{1}=t_{2}+t_{3}=0)$.
@@
(of Lemma 3.8) The proof is elementary and can be checked directly by the definition of the Voisin map $\varphi$. For example to compute $\varphi(E_{02},E_{13})$, consider the plane $\pi=(t_{1}=0)$ containing the line $E_{02}$ and the point $p\in E_{13}$. Its intersection with $S$ is given by $\pi\cap S=E_{02}\cup E_{03}\cup E_{23}$, hence $\varphi(E_{02},E_{13})=\left[E_{03}\cup E_{13}\cup E_{23}\right]$ and the claim is shown.
@@
Let $S$ be a cubic surface with a singular point $P_{2}$ of type $A_{3}$, and 2 singular points $P_{0},P_{3}$ of type $A_{1}$. It contains exactly 5 lines as depicted in Figure 1(b): the line $R_{23}$ through $P_{2},P_{3}$; the line $R_{02}$ through $P_{0},P_{2}$; the line $R_{03}$ through $P_{0}$, $P_{3}$; the line $L_{1}$ containing the only singular point $P_{2}$, and the line $L_{2}$ contained in the smooth locus.
@@
The lines are then described by the equations $R_{23}=(t_{0}=t_{1}=0),R_{02}=(t_{1}=t_{3}=0)$ and $R_{03}=(t_{1}=t_{2}=0)$, $L_{1}=(t_{0}=t_{3}=0)$, and $L_{2}=(t_{2}=t_{0}+t_{3}=0)$.
@@
(of Proposition 3.10) The proof can be checked through straightforward computations. For example, intersecting the surface $S$ with the plane $\pi=(t_{0}=0)$, one gets:
@@
whose singular locus consists of the 2 points $P_{1}:=(0:0:0:1)$ of type $A_{5}$ and $P_{2}:=(0:0:1:0)$ of type $A_{1}$. There are exactly 2 lines on $S$ and they have equations $L_{1}=(t_{0}=t_{1}=0)$ and $L_{2}=(t_{0}=t_{3}=0)$.
@@
(of Proposition 3.12) This follows from the straightforward computation
$(t_{3}=0)\cap S=3L_{2},\quad(t_{0}=0)\cap S=L_{1}+2L_{2}.$
We now study non-normal cubic surfaces, whose analysis is necessary to prove Corollary 3.15.
@@
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$.
Proposition A.1.
@@
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi_{(a:b)}=(at_{0}+bt_{1}=0)$ with $(a:b)\in\mathbb{P}^{1}$. By considering the intersection with such a plane $\pi_{(-1:b)}$ one has:
$X_{6}\cap(-t_{0}+bt_{1}=0)\ =\ \Bigl(-t_{0}+bt_{1}=t_{1}^{2}(b^{2}t_{2}+t_{3})=0\Bigr)\ =\ 2L+R_{b}$
with $R_{b}=\Bigl(b^{2}t_{2}+t_{3}=bt_{1}+t_{0}=0\Bigr)$ different from $L$, and
$X_{6}\cap(t_{1}=0)\ =\ \Bigl(t_{1}=t_{0}^{2}t_{2}=0\Bigr)\ =\ 2L+R$
@@
$t_{0}t_{1}t_{2}+t_{0}^{2}t_{3}+t_{1}^{3}=0.$
It is a non-normal surface whose singular locus consists of the line $L=(t_{0}=t_{1}=0)$.
Proposition A.2.
@@
The proof is an elementary computation. Each plane containing $L$ is of the form $\pi_{(a:b)}=(at_{0}+bt_{1}=0)$ with $(a:b)\in\mathbb{P}^{1}$. By considering the intersection with such a plane $\pi_{(a:-1)}$ one has:
$X_{7}\cap(at_{0}-t_{1}=0)\ =\ \Bigl(at_{0}-t_{1}=t_{0}^{2}(a^{3}t_{0}+at_{2}+t_{3})=0\Bigr)\ =\ 2L+R_{a}.$
with $R_{a}=(a^{3}t_{0}+at_{2}+t_{3}=t_{1}-at_{0}=0)$. Also, one has:
$X_{7}\cap(t_{0}=0)\ =\ \Bigl(t_{1}^{3}=t_{0}=0\Bigr)\ =3L.$
Hence $\pi=(t_{0}=0)$ is the only plane as in the claim. ∎
@@
$t_{1}^{3}+t_{2}^{3}+t_{1}t_{2}t_{3}=0.$
It is a non-normal surface isomorphic to the cone over a nodal planar cubic whose singular locus consists of the line $L=(t_{1}=t_{2}=0)$.
Proposition A.3.
@@
The proof is an elementary computation: Each plane containing $L$ is of the form $\pi_{(a:b)}=(at_{1}+bt_{2}=0)$ with $(a:b)\in\mathbb{P}^{1}$. By considering the intersection with such a plane $\pi_{(1:-b)}$ one has:
$X_{8}\cap(t_{1}-bt_{2}=0)\ =\ \Bigl(t_{1}-bt_{2}=t_{2}^{2}((b^{3}+1)t_{2}+bt_{3})=0\Bigr)\ =\ 2L+R_{b}.$
with $R_{b}=(t_{1}-bt_{2}=(b^{3}+1)t_{2}+bt_{3}=0)$. Also, one has:
$X_{8}\cap(t_{2}=0)\ =\ \Bigl(t_{1}^{3}=t_{2}=0\Bigr)\ =\ 3L.$